GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

GNU and Linux commands

About my favourite file manager Linux for Teachers

Nick Urbanik <nicku@nicku.org>

This document Licensed under GPL—see section 8

5, 6 December 2005

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Outline

Why the Command Line?

Advantages of the Command Line Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands?

But where is the woman command? Isn’t this sexist? What’s the GNU, anyway? Some Things to Do

Getting more software Some Things to Read License Of This Document

GNU and Linux commands Nick Urbanik

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Why the Command Line?

▶ Isn’t that going back to the dark ages? ▶ Don’t file managers make all that obsolete? ▶ Doesn’t it take so much longer to type all those crazy

commands? ▶ “Aren’t you just telling us all this crap to waste our precious

time we could spend learning something useful?” ▶ I’ll tell you a secret

GNU and Linux commands Nick Urbanik

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Why the Command Line?

▶ Isn’t that going back to the dark ages? ▶ Don’t file managers make all that obsolete? ▶ Doesn’t it take so much longer to type all those crazy

commands? ▶ “Aren’t you just telling us all this crap to waste our precious

time we could spend learning something useful?” ▶ I’ll tell you a secret

GNU and Linux commands Nick Urbanik

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Why the Command Line?

▶ Isn’t that going back to the dark ages? ▶ Don’t file managers make all that obsolete? ▶ Doesn’t it take so much longer to type all those crazy

commands? ▶ “Aren’t you just telling us all this crap to waste our precious

time we could spend learning something useful?” ▶ I’ll tell you a secret

GNU and Linux commands Nick Urbanik

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Why the Command Line?

▶ Isn’t that going back to the dark ages? ▶ Don’t file managers make all that obsolete? ▶ Doesn’t it take so much longer to type all those crazy

commands? ▶ “Aren’t you just telling us all this crap to waste our precious

time we could spend learning something useful?” ▶ I’ll tell you a secret

GNU and Linux commands Nick Urbanik

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Why the Command Line?

▶ Isn’t that going back to the dark ages? ▶ Don’t file managers make all that obsolete? ▶ Doesn’t it take so much longer to type all those crazy

commands? ▶ “Aren’t you just telling us all this crap to waste our precious

time we could spend learning something useful?” ▶ I’ll tell you a secret

GNU and Linux commands

Nick Urbanik

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands?

What’s the GNU, anyway?

Some Things to Do

Some Things to Read

License Of This Document

My Dirty Secret

▶ You may be better with a file manager than I am ▶ you may have had more practice than I have… ▶ … because my favourite file manager is the command line

▶ … keeps umbrella handy to protect against the expected shower of rotten tomatoes

Nick Urbanik

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands?

What’s the GNU, anyway?

Some Things to Do

Some Things to Read

License Of This Document

My Dirty Secret

▶ You may be better with a file manager than I am ▶ you may have had more practice than I have… ▶ … because my favourite file manager is the command line

▶ … keeps umbrella handy to protect against the expected shower of rotten tomatoes

Nick Urbanik

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands?

What’s the GNU, anyway?

Some Things to Do

Some Things to Read

License Of This Document

My Dirty Secret

▶ You may be better with a file manager than I am ▶ you may have had more practice than I have… ▶ … because my favourite file manager is the command line

▶ … keeps umbrella handy to protect against the expected shower of rotten tomatoes

Nick Urbanik

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands?

What’s the GNU, anyway?

Some Things to Do

Some Things to Read

License Of This Document

My Dirty Secret

▶ You may be better with a file manager than I am ▶ you may have had more practice than I have… ▶ … because my favourite file manager is the command line

▶ … keeps umbrella handy to protect against the expected shower of rotten tomatoes

Nick Urbanik

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands?

What’s the GNU, anyway?

Some Things to Do

Some Things to Read

License Of This Document

My Dirty Secret

▶ You may be better with a file manager than I am ▶ you may have had more practice than I have… ▶ … because my favourite file manager is the command line

▶ … keeps umbrella handy to protect against the expected shower of rotten tomatoes

GNU and Linux commands Nick Urbanik

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Outline

Why the Command Line?

Advantages of the Command Line Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands?

But where is the woman command? Isn’t this sexist? What’s the GNU, anyway? Some Things to Do

Getting more software Some Things to Read License Of This Document

Nick Urbanik

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands?

What’s the GNU, anyway?

Some Things to Do

Some Things to Read

License Of This Document

Advantages of the Command Line

▶

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Comparing GNU commands with DOS

▶ You are nearly as old as I am, ▶ so many of you are familiar with DOS commands. ▶ If you were just out of school, this comparison would be

useless

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

GNU leftrightarrow DOS

Action GNU  DOS
cp ⟨sourcefile⟩ ...⟨target⟩ copy ⟨sourcefile⟩ ⟨target⟩
mv ⟨sourcefile⟩ ...⟨target⟩ move ⟨sourcefile⟩ ⟨target⟩
cd ⟨dirname⟩  cd ⟨dirname⟩
mkdir ⟨dirname⟩ ...        md ⟨dirname⟩
rmdir ⟨dirname⟩ ...        rd ⟨dirname⟩

pwd                       cd

cat ⟨file⟩  type ⟨file⟩

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Isn’t it too slow?

▶ The

Nick Urbanik

Why the Command Line?

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands?

But where is the woman command? Isn’t this sexist?

What’s the GNU, anyway?

Some Things to Do

Some Things to Read

License Of This Document

How can I get help with these commands?

▶

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands?

But where is the woman command? Isn’t this sexist?

What’s the GNU,

anyway?

Some Things to Do Some Things to Read License Of This Document

Outline

Why the Command Line?

Advantages of the Command Line Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands?

But where is the woman command? Isn’t this sexist? What’s the GNU, anyway? Some Things to Do

Getting more software Some Things to Read License Of This Document

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

What’s the GNU, anyway?

▶

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

▶

Nick Urbanik

Why the Command Line?

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands?

What’s the GNU, anyway?

Some Things to Do

Getting more software

Some Things to Read

License Of This Document

▶

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do

Getting more software

Some Things to Read License Of This Document

Outline

Why the Command Line?

Advantages of the Command Line Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands?

But where is the woman command? Isn’t this sexist? What’s the GNU, anyway? Some Things to Do

Getting more software Some Things to Read License Of This Document

Nick Urbanik

Why the Command Line?

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands?

What’s the GNU, anyway?

Some Things to Do

Getting more software

Some Things to Read

License Of This Document

▶

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Resources I

Dr. Peter Salus. The Daemon, the GNU & the Penguin. Grocklaw 2005.

http://www.groklaw.net/staticpages/index.php?page=20051013231901859  .

GNU General Public License.

http://www.gnu.org/copyleft/gpl.html

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Topics Covered

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands? But where is the woman command? Isn’t this sexist?

What’s the GNU, anyway?

Some Things to Do Getting more software

Some Things to Read

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Topics Covered

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands? But where is the woman command? Isn’t this sexist?

What’s the GNU, anyway?

Some Things to Do Getting more software

Some Things to Read

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Topics Covered

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands? But where is the woman command? Isn’t this sexist?

What’s the GNU, anyway?

Some Things to Do Getting more software

Some Things to Read

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Topics Covered

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands? But where is the woman command? Isn’t this sexist?

What’s the GNU, anyway?

Some Things to Do Getting more software

Some Things to Read

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Topics Covered

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands? But where is the woman command? Isn’t this sexist?

What’s the GNU, anyway?

Some Things to Do Getting more software

Some Things to Read

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Topics Covered

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands? But where is the woman command? Isn’t this sexist?

What’s the GNU, anyway?

Some Things to Do Getting more software

Some Things to Read

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Topics Covered

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands? But where is the woman command? Isn’t this sexist?

What’s the GNU, anyway?

Some Things to Do Getting more software

Some Things to Read

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Topics Covered

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands? But where is the woman command? Isn’t this sexist?

What’s the GNU, anyway?

Some Things to Do Getting more software

Some Things to Read

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Topics Covered

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands? But where is the woman command? Isn’t this sexist?

What’s the GNU, anyway?

Some Things to Do Getting more software

Some Things to Read

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Topics Covered

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands? But where is the woman command? Isn’t this sexist?

What’s the GNU, anyway?

Some Things to Do Getting more software

Some Things to Read

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

Topics Covered

Why the Command Line?

Advantages of the Command Line

Comparing GNU commands with DOS

Isn’t it too slow?

How can I get help with these commands? But where is the woman command? Isn’t this sexist?

What’s the GNU, anyway?

Some Things to Do Getting more software

Some Things to Read

GNU and Linux commands Nick Urbanik

Why the Command Line? Comparing GNU commands with DOS Isn’t it too slow? How can I get help with these commands? What’s the GNU, anyway? Some Things to Do Some Things to Read License Of This Document

License Of This Document

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2005 Nick Urbanik <nicku@nicku.org> You can redistribute modified or unmodified copies of this document provided that this copyright notice and this permission notice are preserved on all copies under the terms of the GNU General Public License as published by the Free Software Foundation—either version 2 of the License or (at your option) any later version.